# Midichlorian-assisted cold fusion Confidence–Diversity Calibration of AI Judgement Enables Reliable Qualitative Coding

### v1

@@ Abstract
LLMs enable qualitative coding at large scale, but assessing the reliability of their output remains challenging in domains where human experts seldom agree. Analysing 5 680 coding decisions from eight state-of-the-art LLMs across ten thematic categories, we confirm that a model’s mean self-confidence already tracks inter-model agreement closely (Pearson ). Adding model diversity—quantified as the normalised Shannon entropy of the panel’s votes—turns this single cue into a dual signal that explains agreement almost completely (). The confidence–diversity duo enables a three-tier workflow that auto-accepts 35 % of segments with <5 % audit-detected error and routes the remainder for targeted human review, cutting manual effort by up to 65 %. Cross-domain replication on six public datasets spanning finance, medicine, law and multilingual tasks confirms these gains ( improvements of 0.20–0.78). Our results establish a generalisable, evidence-based criterion for calibrating AI judgement in qualitative research.
@@ Introduction
Qualitative research has long faced a structural trade–off between scale and reliability. Traditional workflows typically depend on two–coder adjudication, yet resulting Cohen’s often remains only “moderate” while labour costs grow super–linearly with corpus size.1 Yet in domains lacking a single ground truth, researchers still need principled criteria for deciding when an AI label can be trusted. Drawing on calibration research in metacognitive monitoring, we treat self–confidence as a first–order cue and complement it with panel diversity, an error signal that reveals collective blind spots.2 From a human–behaviour perspective, our study therefore centres on how analysts allocate attention within human–AI collaboration, translating cognitive signals into workflow decisions that optimise expert effort. Computational text-analysis methods—ranging from early topic models to state-of-the-art large language models (LLMs)—are rapidly entering the methodological toolkit of social scientists and health researchers, promising to automate labour-intensive qualitative tasks such as open coding and theme extraction 3, 4, 5, 6, 7, 8. By partitioning the confidence–diversity space into actionable zones, we created a tiered review system that efficiently allocates human attention.
@@
Here we propose to complement confidence with model diversity—the dispersion of predictions across differently trained or prompted LLMs. Diversity has long served as an error bound in ensemble learning14, yet remains under-explored in qualitative research. We analyse eight state-of-the-art LLM variants—four single-response models and their chain-of-thought counterparts—that each coded 71 interview segments against a ten-category framework, yielding 5 680 decisions. A ninth model is introduced only in cross-domain validation (see Methods). Our approach focuses on relatively straightforward coding tasks with conceptually accessible categories that are interpretable to human coders, establishing a foundation for reliable quality assessment. We show that mean self-confidence explains much of the observed inter-model agreement (Pearson ), and that adding model diversity produces a dual indicator that predicts agreement almost perfectly ().
Finally, we show that this confidence–diversity principle generalises across six public datasets spanning finance, medicine and law, and that its predictive accuracy remains robust to the exact weighting of the two signals: although an in-sample grid search favours , the intuitive 0.6/0.4 split performs within 0.03 MAE across domains.
@@
Reliability in qualitative coding. Classic content–analysis texts demonstrate that interpretive categories must be both conceptually coherent and empirically reliable. Formal checks usually rely on chance–corrected agreement metrics such as Cohen’s or Krippendorff’s ; empirical studies and methodological guidelines consistently classify values above as “almost perfect” agreement—and report that such cases are uncommon in practice.15, 16, 11, 17, 18 Recent work therefore explores whether computational models—from earlier topic-modelling pipelines to contemporary LLMs—can assist human coders at scale while preserving rigour, reporting promising early results in political science, communication and social research.6, 3, 4 Yet even with multiple annotators, majority voting improves stability only marginally because humans (and often models) share similar blind spots.
@@
Ensemble voting and uncertainty in NLP. Disagreement among independently prompted or fine–tuned models has been leveraged to flag hard examples, measure trust scores, and reduce hallucinations.23 Recent work on uncertainty–aware majority voting extends this logic to crowdsourced labels.24 Diversity within the ensemble therefore serves as an additional uncertainty cue that complements confidence; classic machine–learning theory offers quantitative diversity indices.25 Our study unifies confidence and diversity in the absence of ground truth, showing that their linear combination predicts agreement with , and applies this mechanism to qualitative coding for the first time.
@@ Results
To contextualise the dual–signal gains we first benchmarked model performance against a traditional two–coder baseline and our full three–tier review workflow (Table 1). Across the 5 680 coding decisions the eight–model majority already surpasses typical human reliability, achieving a chance–corrected Cohen’s and halving the error rate relative to double human coding. Integrating the dual–signal triage and expert reconciliation lifts effective reliability to while trimming the residual error rate to 4.1%. These figures derive from the audited interview corpus of 71 segments (see Methods).
@@
These residuals correlate strongly with model diversity (), defined as the normalised entropy of the eight binary votes. Categories with greater diversity are under-predicted by the confidence-only model, whereas near-consensus categories are over-predicted (Fig. 1B; ). Combining and yields a dual-signal regression plane that captures almost all variance (; Fig. 1C), eliminating bias without additional features. While item‐level cross-validation identifies a pure‐diversity optimum (; Methods, Extended Data Fig. 7), we deliberately adopt the intuitive split throughout the paper because it transfers more reliably to six external datasets (median MAE) and aids substantive interpretation. Representative high-diversity examples are provided in Supplementary Table 5.
@@
At the category level this two-variable model explains of the variance—an improvement of 12 percentage points over the confidence-only baseline (). The mean absolute error drops from 3.06 % to 1.48 %, confirming that diversity corrects systematic biases left by confidence alone. Full coefficient estimates with classical t–based 95 % confidence intervals are reported in Supplementary Table 10; percentile bootstrap intervals based on 1,000 resamples (Supplementary Table 11) show the same directional pattern, confirming that the diversity term remains significantly negative despite the small sample size. An item-level analysis over all 710 coding decisions (Supplementary Table 12) produces virtually identical slopes under both OLS and mixed–effects specifications, providing further evidence of model robustness across levels of aggregation.
@@
All improvements were statistically significant at using 5,000 bootstrap resamples, confirming that these substantial gains are not attributable to chance. The consistent performance across this diverse range of tasks—from subjective moral reasoning to technical legal classification—underscores the generality of the confidence–diversity principle for calibrating AI judgements in qualitative analysis.
@@ Discussion
Our findings establish three headline results. (i) A simple dual indicator—mean confidence plus vote diversity—explains 98% of variance in inter-model agreement, doubling the accuracy of confidence alone. (ii) Routing coding points by this indicator enables a three-tier workflow that auto-accepts one-third of decisions and cuts human effort by 65% while preserving accuracy. (iii) Both expert adjudication and tests on six external datasets confirm that these gains generalise beyond the original interview corpus. The distribution of errors across risk tiers represents a particularly important finding: 87% of all errors concentrate in just 50% of the data (the high-risk tier), while the low-risk tier accounts for 35% of data but only 5% of errors. This asymmetric error concentration is what makes our targeted approach so efficient. Similarly, the near-perfect predictive power of the dual-signal model (reducing mean absolute error by half compared to confidence-only) provides a reliable foundation for implementing quality control mechanisms in qualitative research at scale.
@@
The dual-signal calibration approach represents a novel paradigm for model calibration in qualitative research. While previous work has primarily relied on single metacognitive signals such as confidence, our integration of model diversity provides a more robust framework for reliability assessment. This builds upon theoretical foundations in ensemble learning, where diversity has long served as an error bound, but extends this concept to qualitative coding in a novel application. The optimized risk score formula balances these signals based on empirical evidence, creating a transparent and adjustable mechanism for risk assessment. Our emphasis on diversity resonates with deep‐ensemble strategies … 21, 29, 25 and complements post-hoc reliability metrics such as the trust score proposed by Jiang et al.23, thereby bridging machine-learning calibration research with qualitative methodology.
@@
The high predictive capacity of our dual-signal model () allows researchers to reliably forecast inter-model agreement. In practical terms, qualitative researchers can quantitatively estimate coding reliability without extensive human validation or ground truth labels—a significant advancement for resource-constrained projects. This 50% MAE reduction roughly halves misclassification and can save many hours of manual correction on large projects.
@@
Several limitations warrant consideration. First, diversity measurement has reduced resolution in smaller panels (e.g., four‐model panels), limiting the granularity of risk assessment14 (see Supplementary Table 6 and Extended Data Fig. 6). Second, the dual–signal regression in Fig. 1 is fitted on ten category–level observations, providing an illustrative but low-powered test (). Bootstrap validation (Supplementary Table 11) confirms the robustness of the diversity coefficient, yet the wide intervals for the intercept and confidence term reflect this sample constraint. Nevertheless, the six-dataset cross-domain validation (Results; Extended Data Fig. 7) reproduces the same directional pattern, substantially mitigating concerns about limited statistical power. An item-level mixed-effects re-analysis (see Supplementary Table 12) yields similar slopes but narrower intervals, supporting our qualitative conclusions. Third, for the binary Finance–News benchmark the baseline agreement is already near ceiling (83 % unanimous labels), which mechanically caps explainable variance such that the dual-signal model raises only marginally (0.1460.154). We interpret this as a ceiling effect rather than a failure of the method. Fourth, our approach is validated on relatively straightforward coding tasks with conceptually accessible categories. More complex, nuanced coding scenarios—such as highly abstract theoretical constructs, deeply contextual interpretations, or tasks requiring extensive domain expertise—may require different calibration approaches and warrant separate investigation. Fifth, conceptual clarity significantly impacts method effectiveness, with abstract categories requiring more human intervention regardless of confidence levels. Finally, language and cultural specificity may present challenges for cross-lingual applications that require further investigation.
@@
Future work should extend this approach to additional domains and task types, including multi-label classification, sequence labeling, cross-cultural qualitative analysis, and more complex interpretive frameworks.30 A particularly important direction involves investigating how confidence-diversity calibration performs on highly complex coding tasks that require deep contextual understanding, abstract theoretical reasoning, or extensive domain expertise. Such tasks may necessitate modified calibration approaches, additional metacognitive signals, or hybrid human-AI workflows that differ from our current three-tier system. Refinements to the risk scoring formula could incorporate additional metacognitive signals or adjust weights based on domain-specific requirements. Dynamic threshold systems that automatically adjust based on real-time feedback represent another promising direction. Additionally, developing specialized expert-AI interaction interfaces could further optimize human-machine collaborative efficiency.
@@ Conclusion
The dual-signal calibration method provides a reliable framework for AI-assisted coding in qualitative research, while the three-tier workflow significantly improves efficiency without compromising quality. The cross-domain applicability offers researchers across disciplines a practical tool for enhancing their methodological toolkit. Human-AI collaborative coding represents an important methodological advancement that respects the interpretive nature of qualitative research while addressing its scalability challenges.
@@ Methods
For each model–segment–category triple we recorded the binary coding decision, a self-reported confidence score and, when permitted, the full chain-of-thought trace. Reliability was characterised by two orthogonal signals: (i) the mean confidence across the eight models; (ii) decision diversity , computed as the normalised Shannon entropy of the vote distribution. These quantities were combined into a risk score
@@
A stratified random sample of coding points from each review tier (total ) was independently annotated by three trained qualitative researchers who were blinded to model identity (Cohen’s ). Items lacking a two-coder majority proceeded to expert adjudication by two senior scholars in qualitative methodology. The adjudicated label served as ground truth for evaluating accuracy, precision, recall and scores. The green-zone residual-error rate reported in the Results section is derived from this audit: it equals the fraction of audited auto-accepted points whose AI label disagreed with expert consensus, extrapolated to the full green zone.
@@
All analyses were conducted in Python 3.10 using pandas 2.2, statsmodels 0.14 and SciPy 1.12, and correlation coefficients are reported as Pearson’s 37. Ninety-five-per-cent confidence intervals were obtained via -fold non-parametric bootstrap over interview segments. Linear and multiple regressions were fit with ordinary least squares; predictive performance was assessed with nested ten-fold cross-validation using mean absolute error (MAE) and . Correlation coefficients are reported as Pearson’s . Hypothesis tests were two-sided with , and we controlled the false discovery rate using Benjamini–Hochberg correction. Power analysis conducted in G*Power 3.1 indicated that the human-validation sample affords power to detect a percentage-point difference in accuracy between tiers.
@@
All analyses rely on the ten-theme coding scheme originally developed for the “Academic Identity and Knowledge Strategies” project. Full operational definitions and exemplar quotes are provided in Supplementary Table 2; the plain-language label set is archived with the code repository. During inference each LLM received the category name, its two-sentence definition, and an abridged decision rule. Prompts and exact JSON payloads are included in the public release (see Code availability).

### v2

# Midichlorian-assisted cold fusion A Confidence–Diversity Framework for Calibrating AI Judgement in Accessible Qualitative Coding Tasks

@@ Abstract
LLMs enable qualitative coding at large scale, but assessing reliability remains challenging where human experts seldom agree. We investigate confidence-diversity calibration as a quality assessment framework for accessible coding tasks where LLMs already demonstrate strong performance but exhibit overconfidence. Analysing 5 680 coding decisions from eight state-of-the-art LLMs across ten categories, we find that mean self-confidence tracks inter-model agreement closely (Pearson ). Adding model diversity—quantified as normalised Shannon entropy—produces a dual signal explaining agreement almost completely (), though this high predictive power likely reflects task simplicity for current LLMs. The framework enables a three-tier workflow auto-accepting 35 % of segments with <5 % error, cutting manual effort by 65 %. Cross-domain validation confirms transferability ( improvements of 0.20–0.78). While establishing a methodological foundation for AI judgement calibration, the true potential likely lies in more challenging scenarios where LLMs may demonstrate comparative advantages over human cognitive limitations.
@@ Introduction
Qualitative research has long faced a structural trade–off between scale and reliability. Traditional workflows typically depend on two–coder adjudication, yet resulting Cohen’s often remains only “moderate” while labour costs grow super–linearly with corpus size.1 Yet in domains lacking a single ground truth, researchers still need principled criteria for deciding when an AI label can be trusted. We investigate this challenge in accessible coding tasks where LLMs already demonstrate strong performance but exhibit overconfidence. Drawing on calibration research in metacognitive monitoring, we treat self–confidence as a first–order cue and complement it with panel diversity, an error signal that reveals collective blind spots.2 Our study centres on how analysts allocate attention within human–AI collaboration, translating cognitive signals into workflow decisions that optimise expert effort. Computational text-analysis methods—ranging from early topic models to state-of-the-art large language models (LLMs)—are rapidly entering the methodological toolkit of social scientists and health researchers, promising to automate labour-intensive qualitative tasks such as open coding and theme extraction 3, 4, 5, 6, 7, 8. By partitioning the confidence–diversity space into actionable zones, we created a tiered review system that efficiently allocates human attention.
@@
Here we propose to complement confidence with model diversity—the dispersion of predictions across differently trained or prompted LLMs. Diversity has long served as an error bound in ensemble learning14, yet remains under-explored in qualitative research. We analyse eight state-of-the-art LLM variants—four single-response models and their chain-of-thought counterparts—that each coded 71 interview segments against a ten-category framework, yielding 5 680 decisions. A ninth model is introduced only in cross-domain validation (see Methods). Our approach focuses on coding tasks that, while presenting moderate challenges for human coders (), are relatively straightforward for LLMs (). This complexity differential provides an ideal testbed for developing calibration methods. We show that mean self-confidence explains much of the observed inter-model agreement (Pearson ), and that adding model diversity produces a dual indicator that predicts agreement almost perfectly ()—though this high predictive power likely reflects task accessibility for current LLMs.
Finally, we show that this confidence–diversity principle transfers across six public datasets spanning finance, medicine and law within similar complexity ranges, and that its predictive accuracy remains robust to the exact weighting of the two signals: although an in-sample grid search favours , the intuitive 0.6/0.4 split performs within 0.03 MAE across domains. However, these validation tasks share the characteristic of being relatively accessible to current LLMs.
@@
Reliability in qualitative coding. Classic content–analysis texts demonstrate that interpretive categories must be both conceptually coherent and empirically reliable. Formal checks usually rely on chance–corrected agreement metrics such as Cohen’s or Krippendorff’s ; empirical studies and methodological guidelines consistently classify values above as “almost perfect” agreement—and report that such cases are uncommon in practice.15, 16, 11, 17, 18 Recent work therefore explores whether computational models—from earlier topic-modelling pipelines to contemporary LLMs—can assist human coders at scale while preserving rigour, reporting promising early results in political science, communication and social research.6, 3, 4 Yet even with multiple annotators, majority voting improves stability only marginally because humans (and often models) share similar blind spots, particularly in accessible coding scenarios where individual models already perform well.
@@
Ensemble voting and uncertainty in NLP. Disagreement among independently prompted or fine–tuned models has been leveraged to flag hard examples, measure trust scores, and reduce hallucinations.23 Recent work on uncertainty–aware majority voting extends this logic to crowdsourced labels.24 Diversity within the ensemble therefore serves as an additional uncertainty cue that complements confidence; classic machine–learning theory offers quantitative diversity indices.25 Our study unifies confidence and diversity in the absence of ground truth for accessible qualitative coding tasks, showing that their linear combination predicts agreement with in this specific domain, establishing a methodological foundation for more challenging scenarios.
@@ Results
To contextualise the dual–signal gains we first benchmarked model performance against a traditional two–coder baseline and our full three–tier review workflow (Table 1). Across the 5 680 coding decisions the eight–model majority already surpasses typical human reliability, achieving a chance–corrected Cohen’s and halving the error rate relative to double human coding. This substantial performance advantage reflects the relative accessibility of our coding tasks for current LLMs. Integrating the dual–signal triage and expert reconciliation lifts effective reliability to while trimming the residual error rate to 4.1%. These figures derive from the audited interview corpus of 71 segments (see Methods).
@@
These residuals correlate strongly with model diversity (), defined as the normalised entropy of the eight binary votes. Categories with greater diversity are under-predicted by the confidence-only model, whereas near-consensus categories are over-predicted (Fig. 1B; ). Combining and yields a dual-signal regression plane that captures almost all variance (; Fig. 1C), eliminating bias without additional features. This exceptionally high predictive power likely reflects the relative straightforwardness of our coding tasks for state-of-the-art LLMs. While item‐level cross-validation identifies a pure‐diversity optimum (; Methods, Extended Data Fig. 7), we deliberately adopt the intuitive split throughout the paper because it transfers more reliably to six external datasets (median MAE) and aids substantive interpretation. Representative high-diversity examples are provided in Supplementary Table 5.
@@
At the category level this two-variable model explains of the variance—an improvement of 12 percentage points over the confidence-only baseline (). The mean absolute error drops from 3.06 % to 1.48 %, confirming that diversity corrects systematic biases left by confidence alone within this accessible coding domain. Full coefficient estimates with classical t–based 95 % confidence intervals are reported in Supplementary Table 10; percentile bootstrap intervals based on 1,000 resamples (Supplementary Table 11) show the same directional pattern, confirming that the diversity term remains significantly negative despite the small sample size. An item-level analysis over all 710 coding decisions (Supplementary Table 12) produces virtually identical slopes under both OLS and mixed–effects specifications, providing further evidence of model robustness across levels of aggregation.
@@
All improvements were statistically significant at using 5,000 bootstrap resamples, confirming that these substantial gains are not attributable to chance. The consistent performance across this diverse range of tasks—from subjective moral reasoning to technical legal classification—underscores the generality of the confidence–diversity principle for calibrating AI judgements in accessible qualitative analysis tasks where LLMs demonstrate strong baseline performance.
@@ Discussion
Our findings establish three headline results within the domain of accessible qualitative coding tasks. (i) A simple dual indicator—mean confidence plus vote diversity—explains 98% of variance in inter-model agreement, doubling the accuracy of confidence alone. (ii) Routing coding points by this indicator enables a three-tier workflow that auto-accepts one-third of decisions and cuts human effort by 65% while preserving accuracy. (iii) Both expert adjudication and tests on six external datasets confirm that these gains generalise across similar complexity ranges. The exceptionally high predictive power likely reflects the relative accessibility of these coding tasks for current LLMs. The distribution of errors across risk tiers represents a particularly important finding: 87% of all errors concentrate in just 50% of the data (the high-risk tier), while the low-risk tier accounts for 35% of data but only 5% of errors. This asymmetric error concentration is what makes our targeted approach so efficient. Similarly, the near-perfect predictive power of the dual-signal model (reducing mean absolute error by half compared to confidence-only) provides a reliable foundation for implementing quality control mechanisms in qualitative research at scale.
@@
The dual-signal calibration approach represents a novel paradigm for model calibration in accessible qualitative research tasks. While previous work has primarily relied on single metacognitive signals such as confidence, our integration of model diversity provides a more robust framework for reliability assessment within this domain. This builds upon theoretical foundations in ensemble learning, where diversity has long served as an error bound, but extends this concept to qualitative coding where LLMs already demonstrate strong baseline performance. The optimized risk score formula balances these signals based on empirical evidence, creating a transparent and adjustable mechanism for risk assessment in scenarios where confidence and diversity signals are particularly informative. Our emphasis on diversity resonates with deep‐ensemble strategies … 21, 29, 25 and complements post-hoc reliability metrics such as LIME23, thereby bridging machine-learning calibration research with qualitative methodology.
@@
The high predictive capacity of our dual-signal model () allows researchers to reliably forecast inter-model agreement within accessible coding domains. This exceptional predictive power, while likely reflecting the relative straightforwardness of our tasks for current LLMs, provides a methodological foundation that can be extended to more challenging scenarios. In practical terms, qualitative researchers can quantitatively estimate coding reliability without extensive human validation or ground truth labels—a significant advancement for resource-constrained projects working with similar task complexities.
@@
Several limitations warrant consideration. First, diversity measurement has reduced resolution in smaller panels (e.g., four‐model panels), limiting the granularity of risk assessment14 (see Supplementary Table 6 and Extended Data Fig. 6). Second, the dual–signal regression in Fig. 1 is fitted on ten category–level observations, providing an illustrative but low-powered test (). Bootstrap validation (Supplementary Table 11) confirms the robustness of the diversity coefficient, yet the wide intervals for the intercept and confidence term reflect this sample constraint. Nevertheless, the six-dataset cross-domain validation (Results; Extended Data Fig. 7) reproduces the same directional pattern, substantially mitigating concerns about limited statistical power. An item-level mixed-effects re-analysis (see Supplementary Table 12) yields similar slopes but narrower intervals, supporting our qualitative conclusions. Third, for the binary Finance–News benchmark the baseline agreement is already near ceiling (83 % unanimous labels), which mechanically caps explainable variance such that the dual-signal model raises only marginally (0.1460.154). We interpret this as a ceiling effect rather than a failure of the method. Fourth, our approach is validated on relatively straightforward coding tasks with conceptually accessible categories. A fundamental limitation is that our validation occurs in a domain where LLMs already outperform human coders ( vs ), potentially inflating the apparent effectiveness of our calibration approach. The exceptional R² values observed may not generalize to more complex interpretive tasks where the confidence-diversity relationship could be fundamentally different. More complex, nuanced coding scenarios—such as highly abstract theoretical constructs, deeply contextual interpretations, or tasks requiring extensive domain expertise—may require different calibration approaches and warrant separate investigation. Paradoxically, the true potential of LLM-assisted qualitative coding may lie in ultra-complex scenarios involving thousands of tokens where human cognitive limitations become more pronounced while LLMs can maintain consistency across vast textual spans.
@@
Future work should prioritize validation on ultra-complex coding scenarios where LLMs may demonstrate their greatest comparative advantages: processing vast textual corpora, maintaining consistency across thousands of tokens, and identifying subtle patterns that exceed human working memory limitations. Such investigations will reveal whether confidence-diversity calibration scales to tasks where AI assistance transitions from convenience to necessity. A particularly important direction involves investigating how our framework performs when the current performance hierarchy (LLM ¿ human) is reversed in truly challenging interpretive tasks. Additional domains for exploration include multi-label classification, sequence labeling, cross-cultural qualitative analysis, and more complex interpretive frameworks that push the boundaries of current LLM capabilities.30 Refinements to the risk scoring formula could incorporate additional metacognitive signals or adjust weights based on domain-specific requirements.
@@ Conclusion
The dual-signal calibration method provides a reliable framework for AI-assisted coding in accessible qualitative research tasks, while the three-tier workflow significantly improves efficiency without compromising quality within this domain. The cross-domain applicability across similar complexity ranges offers researchers a practical tool for enhancing their methodological toolkit. While our approach establishes a methodological foundation for quality assessment, the true potential of LLM-assisted qualitative coding likely lies in more challenging scenarios where LLMs may demonstrate comparative advantages over human cognitive limitations. Human-AI collaborative coding represents an important methodological advancement that respects the interpretive nature of qualitative research while addressing its scalability challenges.
@@ Methods
For each model–segment–category triple we recorded the binary coding decision, a self-reported confidence score and, when permitted, the full chain-of-thought trace. Reliability was characterised by two orthogonal signals: (i) the mean confidence across the eight models; (ii) decision diversity , computed as the normalised Shannon entropy of the vote distribution. This dual-signal approach proved particularly effective for accessible coding tasks where LLMs demonstrate strong baseline performance. These quantities were combined into a risk score
@@
A stratified random sample of coding points from each review tier (total ) was independently annotated by three trained qualitative researchers who were blinded to model identity (Cohen’s ). Items lacking a two-coder majority proceeded to expert adjudication by two senior scholars in qualitative methodology. The adjudicated label served as ground truth for evaluating accuracy, precision, recall and scores. Notably, the human validation revealed that LLMs achieved higher reliability than traditional two-coder approaches on these accessible coding tasks, consistent with our finding that task complexity relative to model capabilities influences calibration effectiveness. The green-zone residual-error rate reported in the Results section is derived from this audit: it equals the fraction of audited auto-accepted points whose AI label disagreed with expert consensus, extrapolated to the full green zone.
@@
All analyses were conducted in Python 3.10 using pandas 2.2, statsmodels 0.14 and SciPy 1.12, and correlation coefficients are reported as Pearson’s 37. Ninety-five-per-cent confidence intervals were obtained via -fold non-parametric bootstrap over interview segments. Linear and multiple regressions were fit with ordinary least squares; predictive performance was assessed with nested ten-fold cross-validation using mean absolute error (MAE) and . Hypothesis tests were two-sided with , and we controlled the false discovery rate using Benjamini–Hochberg correction. Power analysis conducted in G*Power 3.1 indicated that the human-validation sample affords power to detect a percentage-point difference in accuracy between tiers.
@@
All analyses rely on the ten-theme coding scheme originally developed for the “Academic Identity and Knowledge Strategies” project. Full operational definitions and exemplar quotes are provided in Supplementary Table 2; the plain-language label set is archived with the code repository. The coding framework features conceptually accessible categories that are well-suited to current LLM capabilities, contributing to the high predictive performance observed in our calibration approach. During inference each LLM received the category name, its two-sentence definition, and an abridged decision rule. Prompts and exact JSON payloads are included in the public release (see Code availability).